### v1

# Predicting the suitability of photocatalysts for water splitting using Koopmans spectral functionals: The case of TiO2 polymorphs

Abstract
Photocatalytic water splitting has attracted considerable attention for renewable energy production. Since the first reported photocatalytic water splitting by titanium dioxide, this material remains one of the most promising photocatalysts, due to its suitable band gap and band-edge positions. However, predicting both of these properties is a challenging task for existing computational methods. Here we show how Koopmans spectral functionals can accurately predict the band structure and level alignment of rutile, anatase, and brookite TiO2 using a computationally efficient workflow that only requires (a) a DFT calculation of the photocatalyst/vacuum interface and (b) a Koopmans spectral functional calculation of the bulk photocatalyst. The success of this approach for TiO2 suggests that this strategy could be deployed for assessing the suitability of novel photocatalyst candidates.
I Introduction
One of the most pressing problems that we are currently facing is finding easy and low-cost renewable energy sources. Hydrogen production from water is one attractive option. In this process, water is decomposed by visible light into oxygen and hydrogen without the application of external potentials, as was first demonstrated by Fujishima and Honda using $\mathrm{TiO_{2}}$ as an electrode [1]. Ever since, photocatalytic water splitting (PWS) has been in the spotlight as a way to produce hydrogen via renewable energy. Alongside experiments that have explored PWS at a fundamental level — from photon absorption to the production of molecular hydrogen — computational methods have aided our understanding of this process, and are particularly useful for identifying new candidate materials for catalysts. The search for an ideal photocatalytic material is still ongoing, and numerous studies have been conducted to tackle this problem, especially for semiconductor materials (see [2, 3, 4, 5, 6, 7] and references therein).
To better understand the desirable properties for a candidate photocatalyst, let us first briefly revise the PWS process. Water splitting is, of course, not a spontaneous reaction: it requires an energy barrier to be overcome. One of the ways to do this is via solar irradiation. Upon light absorption, the catalyst plays a crucial role in a three-step process: 1. charge carrier generation, in which electrons are promoted from the valence to the conduction band (creating a hole); 2. charge separation of the electron-hole pair and migration to the surface of a catalyst; and 3. participation in redox reactions. The first two steps are especially dependent on the structural and electronic properties of the photocatalytic material. The overall process is given by:
[…]
$\mathrm{4H^{+}+4e^{-}\longrightarrow 2H_{2}(g)\hskip 14.22636ptE_{re}^{o}=0.00V}$ (2)
The water-splitting reaction involves the oxidation of water and reduction of hydrogennder standard conditions the value of the redox potentials — as referenced to the normal hydrogen electrode — are 1.23 V and 0 V respectively [8]. Given these potentials, the band gap of a suitable catalyst needs to be at least 1.23 eV; otherwise the electrons will not have enough energy to start the reaction. In practice, at least 1.6 to 1.8 eV are required, as a certain amount of excess energy (i.e. the so-called “overpotential”) is needed to overcome kinetic barriers and induce the hydrogen and oxygen evolution reactions on the surface of the electrode [9, 10, 11]. The band gap should not be too high either, since that would reduce the amount of visible light that the photocatalyst can harness.
However, an optimal band gap is not enough. For the PWS reaction to occur, the valence band maximum (VBM) needs to be higher than the oxidation potential of water while the conduction band minimum (CBM) needs to be lower than the hydrogen reduction potential. This ensures that the overall reaction has a negative change in Gibbs free energy, and is therefore spontaneous.
Refer to caption Refer to caption Refer to caption
(a) rutile (b) anatase (c) brookite
6 atoms/cell 6 atoms/cell 24 atoms/cell
primitive tetragonal body-centered tetragonal primitive orthorhombic
$P4_{2}/mnm$ $I4_{1}/amd$ $Pbca$
$a=b=4.6$, $c=2.96$ $a=b=3.80$, $c=9.68$ $a=5.14$, $b=9.25$, $c=5.50$
Figure 1: Crystal structures of three $\mathrm{TiO_{2}}$ polymorphs
Given the importance of the band gap and the band alignment on the performance of photocatalysts, we would like to be able to computationally predict these properties. However, this is a challenging task. In computational materials science, Kohn-Sham density functional theory (KS-DFT) is ubiquitous, but its prediction of band gaps and band alignment is unreliable. This failure stems from several shortcomings. Firstly, the Kohn-Sham eigenvalues (from which we obtain both the band gap and band alignment) are actually mathematical construct and do not have a direct physical meaning — they need not correspond to genuine excitation energies of the system [12, 13]. This is true even for the exact exchange-correlation (xc) functional, for which only the highest occupied molecular orbital (HOMO) has an energy that corresponds to a physical excitation. Approximate xc functionals additionally suffer from further errors such as “one-body” self-interaction error (the incomplete cancellation of the Hartree and exchange terms for one-electron systems) [14, 15] and an erroneous curvature of the total energy with respect to the total number of particles (which should in principle be piecewise linear) [16, 17, 18, 19, 20, 21]. The lack of piecewise linearity means that the eigenvalues obtained via approximate xc functionals do not match the total energy differences that one would obtain by explicitly performing electron removal/addition [16, 22, 20].
Higher-order methods have been used to overcome these shortcomings. These include hybrid functionals [23, 24, 16, 25, 26], Hubbard and extended Hubbard functionals [27, 28], and many-body perturbation theories such as GW [29, 30, 31, 32]. The first two approaches (partially) address the issue of piecewise linearity, but still do not have eigenvalues that can be formally interpreted as quasiparticle energies. GW, on the other hand, has formally-well-defined eigenvalues (being a theoretical framework centered on the description of quasiparticles) but (a) it is much more computationally demanding and (b) performing self-consistency does not systematically improve the results — for example, self-consistent quasiparticle GW typically overestimates band gaps and in this regard is less accurate than G0W0 [33].
In this work we employ Koopmans spectral functionals [20] as an alternative approach that provides accurate spectral properties of materials while being less computationally demanding compared to many-body approaches. As a test case, we apply these functionals to the three most stable polymorphs of TiO2: rutile, anatase, and the less-studied brookite [34]. The crystal structures of these three polymorphs are presented in Figure 1.
The paper is organized as follows: in Section II we explain the theoretical framework behind band alignment and how these calculations need to be adapted for Koopmans functionals. In Section III we describe the computational details and methodology that we use. Finally, in Section IV we present and discuss the electronic structure calculations of the three polymorphs and their band alignment via Koopmans functionals, and compare the results against other existing methods.
II Theoretical framework
As mentioned, the conduction and valence band position with respect to the water redox levels is one of the most important features of a promising photocatalyst. The positions of the band edges depend on — among other factors — the crystal structure, the chemical environment, and the surface structure of the material [35, 36, 37]. Broadly speaking, “band alignment” refers to measuring the relative alignment of the energy bands of two materials at an interface. Often this interface is between two solid materials (such as semiconductors) but it can also refer to a junction between a semiconductor and fluid (such as the case of water splitting via photocatalytic surface). There are several methods for calculating the alignment of bands at an interface, which can be broadly categorized based on the choice of reference level, or, relatedly, what system is modeled (e.g. the bulk alone vs. the explicit interface). The band alignment can be estimated from the bulk structure alone via branch point energies (BPEs) [38]. BPEs are defined as the energy points where bands change character from being predominantly donor-like to acceptor-like. It has been shown that the electronegativity and variation in character of interface-induced gap states are determined by the nearest band edge, which can in turn be used to estimate the band offset [39]. One clear advantage of this method is that it only requires calculations of the bulk, drastically reducing its computational cost, but the method cannot be applied to materials for which the BPEs lie in the conduction band region [40]. At the other end of the spectrum, the hetero-interface can be explicitly modeled. For hetero-interfaces without lattice mismatch, this approach is straightforward because it only requires aligning the valence and conduction bands of semiconductors against the calculated average potentials in the plane parallel to the interfaces [41]. In the case of lattice mismatch, potential deformation of a core state needs to be taken into account [42, 43]. Finally, an intermediate approach is to measure the band offsets with respect to a reference level. This reference level can be the vacuum level, in which one obtains the ionization potential (IP) and electron affinity (EA) of each material [44]. These properties are intrinsically surface properties and must be obtained by considering a slab model. Alternatively, an impurity can be used as a reference level, such as hydrogen [45] or transition metal impurities [46], where band alignment relies on the position of dangling bonds formed between the impurity and semiconductor. Other more approximate approaches exist, such as effective dipole moments, tight-binding schemes, and empirical schemes [47, 48]. All of the above methods are nicely summarized in Ref. 49. In this work, we use the method based on the IP and EA of the material, adapted for use with orbital-density-dependent functionals [50, 51, 52, 53].
II.1 Band alignment with DFT
Before discussing how band alignment calculations with respect to the vacuum are performed for orbital-dependent functional theories, we first briefly review how these calculations are performed for standard functionals.
In order to calculate the ionization potential and electron affinity of a material, we must be able to reference the valence band maximum (VBM) and conduction band maximum (CBM) against the vacuum level. However, the VBM and CBM are extracted from DFT calculations as Kohn-Sham eigenvalues, whose absolute value carry no physical meaning: indeed, the Kohn-Sham potential can be shifted by an arbitrary constant leaving the physical system unchanged but shifting the KS eigenvalues by the same amount. To tether the KS eigenvalues to something meaningful, one valid choice is the potential of the vacuum. To obtain the potential in the vacuum, one considers a slab model, where the bulk catalyst is interfaced with vacuum, and then calculates the change in the average electrostatic potential between the slab and vacuum regions (denoted $\Delta V$). This scheme is illustrated in Figure 2.
The average electrostatic potential is obtained by first calculating a planar average of the three-dimensional electrostatic potential:
[…]
where $S$ is the cross-sectional area of the cell parallel to the interface and for simplicity we have assumed a tetragonal cell with a slab oriented in the $z$ direction. Within the slab region, $\overline{V}(z)$ exhibits oscillations with a periodicity matching that of the ionic cores. These oscillations can be removed by macroscopically averaging the potential:
[…]
using an averaging window of length $L$ that matches the periodicity of the lattice.
Having simulated a slab model and extracted $\Delta V$, the vacuum level for a bulk system is then given by the average electrostatic potential across the entire cell $\langle V\rangle_{\mathrm{bulk}}$ plus the potential difference $\Delta V$ as calculated from the slab calculation. The IP and EA are then given as the (negative of the) VBM and CBM energy levels relative to this vacuum level i.e.
$\mathrm{IP}=\Delta V-\varepsilon_{\mathrm{VBM}}$ (5)
and
$\mathrm{EA}=\Delta V-\varepsilon_{\mathrm{CBM}}$ (6)
We stress that this procedure relies on (a) a sufficiently thick slab, so that the electrostatic potential deep within the slab is bulk-like [54], and (b) a sufficiently wide vacuum region, in order for the electrostatic potential to converge. It is also important to allow for structural relaxation of the slab/vacuum interface as this can significantly affect the potential difference $\Delta V$ [55].
Figure 2: Cartoon of the band alignment procedure. The black line represents the macroscopic average potential $\Delta V$ calculated across the slab for most stable surface facet. The IP and EA correspond to the VBM and CBM relative to the vacuum reference level, which is obtained via alignment of the average potential $V_{\mathrm{bulk}}$ for the slab and bulk systems.
This approach has one notable drawback: semi-local DFT eigenvalues are often not quantitatively (nor even qualitatively) accurate. It is therefore necessary to go beyond DFT to obtain accurate predictions of the valence and conduction band positions.
II.2 Koopmans spectral functionals
So why are KS-DFT eigenvalues unreliable, and what can we do to improve them? To start to answer this question, we note that DFT Kohn-Sham eigenvalues do not match the corresponding energy difference one obtains from explicitly performing electron addition/removal (i.e. a $\Delta$SCF calculation). Contrast this with the exact one-body Green’s function — an object that describes one-particle excitations exactly — whose excitation energies (i.e. poles) are located precisely at points that correspond to addition/removal energies (as can be straightforwardly seen from its spectral representation).
Inspired by this observation, Koopmans functionals are a class of functionals that seek to accurately describe the spectral properties of materials by restoring eigenvalue/total-energy-difference equivalence to DFT [20]. To do so, they impose the so-called “generalized piecewise-linearity” (GPWL) condition, which states that the orbital energies $\varepsilon_{i}$ of orbitals $\varphi_{i}$ should be independent of that orbital’s occupation $\emph{f}_{i}$:
$\varepsilon_{i}=\langle\varphi_{i}|\hat{H}|\varphi_{i}\rangle=\frac{dE}{d\emph{f}_{i}}$ (7)
This is related to the more well-known “piecewise linearity” (PWL) condition i.e. linearity of the total energy with respect to the total number of electrons in the system. Standard density functional approximations are not piecewise-linear, causing the aforementioned discrepancy between total energy differences and eigenvalues.
The general form of a functional that imposes the GPWL condition is:
[…]
As this equation shows, Koopmans functionals take the form of a correction to DFT: The first term on the right-hand side, $E^{\mathrm{DFT}}$, is the energy of a (typically semi-local) “base” functional. The corrective terms then impose GPWL (equation 7) by removing, orbital-by-orbital, all non-linear dependence of the total energy on the orbital occupancies (the first term inside the parentheses) and replacing it with a term that is linear with respect to the orbital occupation $\emph{f}_{i}$ (the second term inside the parentheses) [56, 57, 58, 59]. Here $\eta_{i}$ is some constant. Depending on how one chooses $\eta_{i}$, different flavors of Koopmans can be defined, such as KI and KIPZ [56]. The Koopmans integral correction (KI) sets $\eta_{i}$ in order to guarantee that the orbital energies $\varepsilon_{i}$ are equal to the corresponding $\Delta$SCF total energy difference of the base functional. Explicitly, the KI functional is given by
$E^{\mathrm{KI}}[\{\rho_{i}\}]=E^{\mathrm{DFT}}[\rho]+\sum_{i}\alpha_{i}\Big{[}\Big{(}E_{\mathrm{Hxc}}[\rho-\rho_{i}]-E_{\mathrm{Hxc}}[\rho]\Big{)}+\\ f_{i}\Big{(}E_{\mathrm{Hxc}}[\rho-\rho_{i}+n_{i}]-E_{Hxc}[\rho-\rho_{i}]\Big{)}\Big{]}$ (9)
where $\rho_{i}$ is orbital density at filling $f_{i}$ and $n_{i}$ is the normalized orbital density i.e. $\rho_{i}(\mathbf{r})=f_{i}n_{i}(\mathbf{r})$. The parameters $\alpha_{i}$ are screening parameters, which account for the fact that the derivation of the KI correction only accounts for the explicit dependence of the DFT energy on the orbital occupancies in going from equation 8 to 9. By scaling the strength of the orbital correction via these screening parameters, we account for orbital relaxation post hoc. Crucially, these screening parameters $\alpha_{i}$ are system-specific, and can be computed ab initio — i.e. they are not fitting parameters [60].
The second form of Koopmans functionals, KIPZ, takes the KI functional and adds one further term to the functional, a screened Perdew-Zunger (PZ) self-interaction correction term:
$E^{\mathrm{KIPZ}}=E^{\mathrm{KI}}-\sum_{i}\alpha_{i}f_{i}E_{\mathrm{Hxc}}[n_{i}]$ (10)
which removes one-body self-interaction error and guarantees that the functional is exact for one-electron systems. Because evaluating the KIPZ functional is more expensive than the KI functional, it can be advantageous to evaluate the KIPZ Hamiltonian on the KI ground state. This is referred to as perturbative KIPZ (pKIPZ).
An important feature of Koopmans functionals is that they are not density functionals: the energy depends not only on the total density of the system but also on the densities of individual orbitals, making this an “orbital-density-dependent” functional theory. This comes with several important consequences. In the framework of DFT, the functional is invariant with respect to unitary rotations of the occupied manifold, and this means that the same set of orbitals minimize the total energy and diagonalize the Hamiltonian. However this is not the case for orbital-density dependent functionals, for which the “variational” orbitals minimize the energy while the “canonical” orbitals diagonalize the Hamiltonian constructed at the end of minimization procedure. These two sets of orbitals are related via a unitary transformation.
However it is important to note that emergence of localized variational orbitals is crucial because it allows Koopmans functionals to correctly treat bulk systems [57]. The Koopmans correction is applied to each variational orbital, while the canonical orbitals are interpreted as Dyson orbitals and their eigenvalues as quasiparticle energies, which are shifted by some amount depending on the size of the Koopmans corrections applied to each of their constituent variational orbitals. In the simplest case, where all the variational orbitals are symmetrically equivalent, the Koopmans correction results in a constant shift of the valence and conduction bands, opening the band gap. For systems with non-equivalent variational orbitals the Koopmans correction also affects inter- and intra-band distances, bandwidths, effective masses, etc..
Because Koopmans functionals are orbital-density-dependent and have screening parameters that must be calculated ab initio, the procedure for computing a Koopmans band structure involves several steps as follows:
  • a ground-state DFT calculation to initialize the total density
  • a Wannierization of the DFT Kohn-Sham states in order to obtain maximally-localized Wannier orbitals with which to initialize the variational orbitals
  • calculating the screening parameters $\{\alpha_{i}\}$ via finite differences or density functional perturbation theory [61, 62]
  • a final Koopmans calculation to minimize the energy and construct and diagonalize the Hamiltonian
The accuracy of Koopmans functionals has already been proven across a range of different materials [61, 63, 62, 64]. For more details about Koopmans functionals, we refer the reader to Ref. [65].
II.3 Band alignment with Koopmans functionals
To calculate band alignment with Koopmans functionals, we must adapt the procedure described earlier in Section II.1. Here, we take advantage of a peculiar feature of the KI and pKIPZ functionals: for insulating systems where all the orbital occupancies are either 0 or 1, the ground state KI and pKIPZ energies and densities match those of the base functional. (This can be seen by setting $f_{i}$ to 0 or 1 in equation 9). We stress that this does not mean that the Koopmans correction has no effect, because the derivative of the correction is non-vanishing and thus the eigenvalues of the Koopmans functional will be different to those of the base functional. Nevertheless, this means that the ground-state density of Koopmans functionals match those of the base functional, and when performing band alignment the slab calculations only need to be performed at the level of DFT, because — being a ground-state property — the average electrostatic potential remains unchanged by the Koopmans correction [44, 66]. This substantially reduces the cost of this framework for computing band alignment.
Given this simplification, for Koopmans functionals equations 11 and 12 become:
$\mathrm{IP}=\Delta V^{\mathrm{DFT}}-(\varepsilon_{\mathrm{VBM}}^{\mathrm{DFT}}+\Delta\varepsilon_{\mathrm{VBM}})$ (11)
and
$\mathrm{EA}=\Delta V^{\mathrm{DFT}}-(\varepsilon_{\mathrm{CBM}}^{\mathrm{DFT}}+\Delta\varepsilon_{\mathrm{CBM}})$ (12)
where $\Delta\varepsilon_{\mathrm{VBM}}$ and $\Delta\varepsilon_{\mathrm{VBM}}$ are the shifts in these band edges due to the Koopmans correction. Equivalently, one can calculate the electron affinity (EA) by simply subtracting the value of electronic gap calculated of Koopmans level from the IP:
$\mathrm{EA}=\mathrm{IP}-E_{g}^{\mathrm{Koopmans}}$ (13)
III Method
In this work, we present the band gaps and band alignment of three common polymorphs of TiO2. The crystal structure of these three polymorphs were obtained from the Materials Cloud three-dimensional crystal database (MC3D) [67], which are structures whose geometries have been optimized using semi-local DFT. In the case of anatase and brookite, the optimized lattice parameters are $a$ = $b$ = 3.8001 Å, and $c$ = 9.6814 Å for anatase and $a$ = 5.1689 Å, $b$ = 9.2548 Å, and $c$ = 5.5035 Å for brookite. They differ by less than 2% from experimentally-reported values ($a$ = $b$ = 3.7842 Å and $c$ = 9.5146 Å for anatase and $a$ = 5.13 Å, $b$ = 9.16 Å, and $c$ = 5.43 Å for brookite). For rutile, the optimized lattice parameters are $a$ = 4.60 Å and $c$ = 2.96 Å (which is in excellent agreement with the experimental $a$ = $b$ = 4.5937 Å and $c$ = 2.9581 Å). There are many reported studies, both experimental and theoretical, regarding the crystal structure of TiO2 [68, 69, 70, 71, 72, 73, 74].
Koopmans functionals calculations were performed using Quantum ESPRESSO via the koopmans package [65, 75, 76]. The screening parameters were calculated via finite-differences, which necessitate the use of a supercell and image charge correction schemes to avoid spurious interactions between periodic images [77]. For these calculations, we used a $2\times 2\times 2$ supercell for rutile and anatase and $2\times 1\times 2$ supercell for brookite. To convert these $\Gamma$-only results to a well-sampled primitive cell the eigenvalues are unfolded to the equivalent primitive cell [63]. We used norm-conserving pseudopotentials from the SG15 library (version 1.0) [78]), and a wave-function energy cutoff of 80 Ry as recommended by force convergence data. All of the input and output files can be found on the Materials Cloud Archive [79].
Figure 3: KI@PBE band structures (blue) with respect to the PBE bands (red); the zero reference energy is set to $\varepsilon_{\mathrm{VBM}}^{\mathrm{DFT}}$.
For the slab calculations, we used the thermodynamically most stable facets of $\mathrm{TiO_{2}}$ polymorphs: the (110) surface facet of rutile, (101) of anatase, and (210) of brookite. These surface orientations have been extensively studied in the literature [80, 81, 82, 83, 84, 85, 86, 87, 88] . The slab models were constructed using six and eight periodic units for rutile and anatase, and six periodic units for the larger brookite unit cell. The depth of the vacuum was chosen to be six times the height of the bulk primitive cell (ensuring correspondence in size between primitive and supercell), which corresponds somewhere between 20 and 30 Å depending on the polymorph. These slab sizes were chosen to ensure the macroscopic average potential was converged to within 0.01 eV. The slab calculations used a $3\times 7\times 1$ $k$-point mesh for rutile and anatase and $3\times 2\times 1$ for brookite. These DFT calculations were also performed using Quantum ESPRESSO. The geometries of the slab structures were optimized in order to obtain an accurate physical picture of the surface/vacuum interface.
IV Results and discussion
IV.1 Band structure
In Figure 3 we present the band structures of three polymorphs of TiO2 calculated via the KI functional, and compare them against those calculated with PBE (the underlying base functional). The band gaps of rutile and brookite TiO2 are direct ($\Gamma\rightarrow\Gamma$), while that of anatase is indirect ($Z\rightarrow\Gamma$) [89, 90, 91]. In all three cases, the KI correction shifts both the valence (downwards) and conduction bands (upwards), leading to the band gap widening relative to the PBE result. (The precise values of these shifts can be found in the Supplemental Material.) The resulting band gaps are presented in Table 1, alongside experimental values, as well as those calculated using hybrid functionals and GW. When comparing against experimental band gaps it is important to account for zero-point renormalization — by subtracting the ZPR (as calculated ab initio) from experimental values one obtains a value that can be fairly compared against the computational results obtained for pristine crystal geometries.
Table 1: Electronic band gaps (in eV) of $\mathrm{TiO_{2}}$ polymorphs at different levels of theory.
rutile anatase brookite
PBE $1.73$ $2.21$ $2.42$
HSE06111[92]222[33] 3.39, 3.61 $3.60$ $3.30$
G0W0@PBE11footnotemark: 1 $3.46$ $3.73$ $3.45$
G0W0@HSE0622footnotemark: 2 $4.73$- --
scQPGW@HSE0622footnotemark: 2 $5.18$- --
KI@PBE $3.29$ $3.80$ $4.03$
pKIPZ@PBE $3.38$ $3.90$ $4.14$
exp -- ZPR $3.34-3.41$ $3.34-3.44$-
exp333[93, 94, 95] $3.03-3.06$ $3.10-3.20$ $3.1-3.4$
$-0.337,$
ZPR444[96, 97, 98] $-0.349,$ $-0.238$-
$-0.314$
For rutile and anatase, the electronic band gaps calculated using KI@PBE are within 0.36 eV of experiment (when accounting for ZPR). This level of accuracy is a drastic improvement on that of the base functional PBE, is comparable to that of G0W0@PBE and hybrid functionals, and is better than G0W0@HSE06 and self-consistent quasiparticle GW. For brookite, no ZPR correction was found in the literature and thus comparison with experiment is not possible.
The results with pKIPZ are very similar to that of KI (plots of pKIPZ@PBE band structures can be found in the Supplemental Material). Compared to the KI functional, pKIPZ reports slightly larger band gaps, especially in the case of brookite. Both KI and pKIPZ assign brookite the largest band gap of the three polymorphs, in contrast to HSE06 and G0W0@PBE which assign it the smallest. But without an accurate measurement of the ZPR in brookite and more precise experimental measurements of the band gap, it is difficult to determine which ordering is correct.
IV.2 Band alignment
Table 2: Ionization potentials and electron affinities in eV obtained on KI and pKIPZ level in comparison with experimental studies of single-crystal polymorphs.
rutile (110) anatase (101) brookite (210)
IP EA IP EA IP EA
PBE $7.22$ $5.55$ $7.11$ $4.90$ $7.17$ $4.78$
HSE0600footnotetext: [33]11footnotemark: 1 $8.66$ $4.99$- -- -- --
HSE06 (QM/MM)111[36] $7.83$- $8.3$- -- --
G0W0@PBE11footnotemark: 1 $7.29$ $3.03$- -- -- --
G0W0@HSE0611footnotemark: 1 $7.92$ $3.19$- -- -- --
scQPGW@HSE0611footnotemark: 1 $8.77$ $3.59$- -- -- --
KI@PBE $8.00$ $4.71$ $8.19$ $4.39$ $8.03$ $4.00$
pKIPZ@PBE $7.83$ $4.44$ $8.00$ $4.10$ $7.84$ $3.71$
exp -- ZPR $8.5$ $4.97-5.00$- -- -- --
exp222[80, 99, 100] $8.20$ $5.14-5.17$ 7.96, 8.07, 8.20 -- -- --
ZPR333[101] $-0.3$ $0.17$- -- -- --
The ionization potentials and electron affinities of the $\mathrm{TiO_{2}}$ polymorphs, as calculated via KI and pKIPZ following the method described in Section II.3, are reported in Table 2 alongside values obtained from other higher levels of theory and experiment. Once again it is important to account for ZPR when comparing against experiment. However, ZPR shifts for the VBM and CBM individually have only been reported for the rutile polymorph, calculated using a generalized Fröhlich model [101]. This technique is more approximate than those that were used to calculate the ZPR shifts in the band gaps listed in Table 1. Note that the combining the shifts for the VBM and CBM predicted by the Fröhlich model results in a ZPR reduction of the band gap (0.47 eV) that is up to 50% larger than those listed in Table 1, so these results should be treated with caution.
That caveat aside, for rutile KI@PBE underestimates the shifted experimental IP by 0.5 eV — comparable to the accuracy of G0W0@HSE06. The EA, on the other hand, is predicted by KI@PBE within the range of values reported experimentally — a marked improvement upon the GW results, and matched only by HSE06.
For anatase and brookite, no ZPR results for the IP and EA were found. The KI and pKIPZ result for the IP of anatase are fractionally lower than that given by HSE06 [36].
Figure 4: Band alignment of TiO2 polymorphs using KI and pKIPZ. Blue rectangles represent the IPs of three polymorphs, while red rectangles the EAs. Experimental values are given as red solid lines (from Refs. [80, 99, 57, 100]).
The band alignment diagram represented in Figure 4 shows the IPs and EAs of the three polymorphs compared against the two redox reaction potentials. The H+/H2 potential lies at -4.44 eV [102], while the H2O/O2 potential is 1.23 eV below this (as per Equation 1) [8]. The positions of the valence and conduction bands should straddle these two potentials i.e. the VBM and CBM should not lie within the two dashed lines in Figure 4. According to our calculations, rutile TiO2 appears not to satisfy this condition, which may explain why — despite having a more desirable band gap than anatase and brookite — this polymorph is a slightly worse catalyst, especially compared to anatase [103]. (Of course, an optimal band alignment does not guarantee photocatalytic success: anatase also exhibits higher surface activity and desirable excitonic properties [103, 104].) Brookite, on the other hand, shows some promise as a photocatalyst: the positions of the valence and the conduction bands seem to be favorable, but its fractionally larger band gap would limit the amount of photons that would be absorbed when subjected to sunlight. In this case, one might be able to engineer the band gap.
The KI and pKIPZ results are qualitatively the same, with pKIPZ reporting a slightly more favorable CBM for rutile relative to the hydrogen reduction reaction. Still, with respect to the value of band gap and band edge positions, both KI and pKIPZ predict that anatase appears to be the most promising photocatalyst of the three.
V Conclusions
In conclusion, this investigation demonstrates the accuracy of Koopmans spectral functionals for calculating the band gap, IP, and EAs of TiO2, finding — in agreement with experiment — that anatase is the most promising photocatalyst of the three. The individual predictions of the ionization potentials and electron affinities were either as or more accurate than the results of hybrid functionals and many-body perturbation theory. At the same time, the band alignment calculations presented in this work are notably simpler, given that this framework only requires Koopmans functional calculations on the primitive cell and semi-local DFT calculations for the slab calculations. In the future, Koopmans functionals could be deployed to screen promising photocatalyst candidates that are much less well-studied.
VI Acknowledgments
MS acknowledges support from EXAF (Excellence in Africa) Research Centre as part of the collaboration ”Embedded exact quantum dynamics for photocatalytic water splitting”. This research was supported by the NCCR MARVEL, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 205602).

### v2

@@
 The search for an ideal photocatalytic material is still ongoing, and numerous studies have been conducted to tackle this problem, especially for semiconductor materials (see [61, 86, 71, 104, 121, 35] and references therein).
-To better understand the desirable properties for a candidate photocatalyst, let us first briefly revise the PWS process.
+To better understand the desirable properties for a candidate photocatalyst, let us first briefly revise the PWS process. The process is schematically presented in Figure 1.
+Figure 1: Schematic illustration of photocatalytic water splitting.
+The conduction-band and valence-band regions are shown in green and red, respectively.
+When a photon (yellow line) with energy equal to or greater than the semiconductor band gap is absorbed, an electron is excited from the valence band to the conduction band, leaving a hole behind.
+The electron and hole subsequently participate in the reduction of hydrogen (top right) and oxidation of water (bottom right).
 Water splitting is, of course, not a spontaneous reaction: it requires an energy barrier to be overcome.
@@
 The first two steps are especially dependent on the structural and electronic properties of the photocatalytic material.
-The overall process is given by:
-[…]
+The overall process is given by
+$\begin{split}\mathrm{2H_{2}O(l)\longrightarrow O_{2}(g)+4H^{+}+4e^{-}\hskip 14.22636ptE_{ox}^{o}=1.23V}\end{split}$ (1)
 $\mathrm{4H^{+}+4e^{-}\longrightarrow 2H_{2}(g)\hskip 14.22636ptE_{re}^{o}=0.00V}$ (2)
-The water-splitting reaction involves the oxidation of water and reduction of hydrogennder standard conditions the value of the redox potentials — as referenced to the normal hydrogen electrode — are 1.23 V and 0 V respectively [8].
+The water-splitting reaction involves the oxidation of water and reduction of hydrogen.
+Under standard conditions the value of the redox potentials — as referenced to the normal hydrogen electrode — are 1.23 V and 0 V respectively [96].
 Given these potentials, the band gap of a suitable catalyst needs to be at least 1.23 eV; otherwise the electrons will not have enough energy to start the reaction.
@@
 primitive tetragonal body-centered tetragonal primitive orthorhombic
-$P4_{2}/mnm$ $I4_{1}/amd$ $Pbca$
-$a=b=4.6$, $c=2.96$ $a=b=3.80$, $c=9.68$ $a=5.14$, $b=9.25$, $c=5.50$
+$P4_{2}/mnm$ $I4_{1}/amd$ PbcaPbca
+$a=b=4.6$, $c=2.96$ Å $a=b=3.80$, $c=9.68$ Å $a=5.14$, $b=9.25$, $c=5.50$ Å
 Figure 2: Crystal structures of three $\mathrm{TiO_{2}}$ polymorphs
@@
 II Theoretical framework
-As mentioned, the conduction and valence band position with respect to the water redox levels is one of the most important features of a promising photocatalyst.
+As has already been mentioned, the valence and conduction band positions with respect to the water redox levels are some of the most important features of a promising photocatalyst.
 The positions of the band edges depend on — among other factors — the crystal structure, the chemical environment, and the surface structure of the material [88, 97, 93].
@@
 All of the above methods are nicely summarized in Ref.
-49.
+48.
 In this work, we use the method based on the IP and EA of the material, adapted for use with orbital-density-dependent functionals [55, 122, 100, 117].
@@
 The average electrostatic potential is obtained by first calculating a planar average of the three-dimensional electrostatic potential:
-[…]
+$\overline{V}(z)=\frac{1}{S}\int_{S}dxdyV(r)$ (3)
 where $S$ is the cross-sectional area of the cell parallel to the interface and for simplicity we have assumed a tetragonal cell with a slab oriented in the $z$ direction.
@@
 These oscillations can be removed by macroscopically averaging the potential:
-[…]
+$\overline{\overline{V}}(z)=\frac{1}{L}\int_{z-\frac{L}{2}}^{z+\frac{L}{2}}\overline{V}(z^{\prime})dz^{\prime}$ (4)
 using an averaging window of length $L$ that matches the periodicity of the lattice.
@@
 II.2 Koopmans spectral functionals
-So why are KS-DFT eigenvalues unreliable, and what can we do to improve them?
-To start to answer this question, we note that DFT Kohn-Sham eigenvalues do not match the corresponding energy difference one obtains from explicitly performing electron addition/removal (i.e.
-a $\Delta$SCF calculation).
-Contrast this with the exact one-body Green’s function — an object that describes one-particle excitations exactly — whose excitation energies (i.e.
-poles) are located precisely at points that correspond to addition/removal energies (as can be straightforwardly seen from its spectral representation).
+[…]
 Inspired by this observation, Koopmans functionals are a class of functionals that seek to accurately describe the spectral properties of materials by restoring eigenvalue/total-energy-difference equivalence to DFT [22].
@@
 The general form of a functional that imposes the GPWL condition is:
-[…]
+$E_{\mathrm{Koopmans}}[\rho,{f_{i}}]=E^{\mathrm{DFT}}[\rho]+\sum_{i}\left(-\int_{0}^{f_{i}}\varepsilon_{i}(f)df+f_{i}\eta_{i}\right)$ (8)
 As this equation shows, Koopmans functionals take the form of a correction to DFT: The first term on the right-hand side, $E^{\mathrm{DFT}}$, is the energy of a (typically semi-local) “base” functional.
 The corrective terms then impose GPWL (equation 7) by removing, orbital-by-orbital, all non-linear dependence of the total energy on the orbital occupancies (the first term inside the parentheses) and replacing it with a term that is linear with respect to the orbital occupation $\emph{f}_{i}$ (the second term inside the parentheses) [84, 85, 9, 10].
-Here $\eta_{i}$ is some constant.
-Depending on how one chooses $\eta_{i}$, different flavors of Koopmans can be defined, such as KI and KIPZ [56].
-The Koopmans integral correction (KI) sets $\eta_{i}$ in order to guarantee that the orbital energies $\varepsilon_{i}$ are equal to the corresponding $\Delta$SCF total energy difference of the base functional.
-Explicitly, the KI functional is given by
+Here $\eta_{i}$ is some constant that will be defined shortly.
+An important feature of Koopmans functionals is that they are not density functionals: the energy depends not only on the total density of the system but also on the densities of individual orbitals, making this an “orbital-density-dependent” functional theory.
+This comes with several important consequences.
+In the framework of DFT, the energy is invariant with respect to unitary rotations of the occupied manifold, and this means that the same set of orbitals minimize the total energy and diagonalize the Hamiltonian.
+This is not the case for orbital-density dependent functionals, for which the “variational” orbitals minimize the energy while the “canonical” orbitals diagonalize the Hamiltonian constructed at the end of minimization procedure.
+(Note that this does not imply that the functional is not variational.
+The loss of unitary invariance and the emergence of variational and canonical orbitals is a feature common to all orbital-density dependent functionals including e.g.
+PZ-SIC.)
+[…]
 $E^{\mathrm{KI}}[\{\rho_{i}\}]=E^{\mathrm{DFT}}[\rho]+\sum_{i}\alpha_{i}\Big[\Big(E_{\mathrm{Hxc}}[\rho-\rho_{i}]-E_{\mathrm{Hxc}}[\rho]\Big)+\\ f_{i}\Big(E_{\mathrm{Hxc}}[\rho-\rho_{i}+n_{i}]-E_{Hxc}[\rho-\rho_{i}]\Big)\Big]$ (9)
@@
 By scaling the strength of the orbital correction via these screening parameters, we account for orbital relaxation post hoc.
-Crucially, these screening parameters $\alpha_{i}$ are system-specific, and can be computed ab initio — i.e.
+Crucially, these screening parameters $\alpha_{i}$ are system-specific, and can be computed ab initioi.e.
 they are not fitting parameters [18].
-The second form of Koopmans functionals, KIPZ, takes the KI functional and adds one further term to the functional, a screened Perdew-Zunger (PZ) self-interaction correction term:
+Curiously, for insulating systems where all the orbital occupancies are either 0 or 1, the ground state KI energies and densities match those of the base functional.
+(This can be seen by setting $f_{i}$ to 0 or 1 in equation 9, in which case the entire correction vanishes.
+(We stress that this does not mean that the Koopmans correction has no effect, because the derivative of the correction is non-vanishing and thus the eigenvalues of the Koopmans functional will be different to those of the base functional.) This property of the KI functional allows us to obtain the KI ground-state density directly from a single semi-local DFT calculation, after which only unitary rotations of the occupied variational orbitals are needed to locate the KI minimum.
+This greatly reduces the computational cost of these calculations.
+The second form of Koopmans functionals, KIPZ, adds a screened Perdew-Zunger (PZ) self-interaction correction term to the functional:
 $E^{\mathrm{KIPZ}}=E^{\mathrm{KI}}-\sum_{i}\alpha_{i}f_{i}E_{\mathrm{Hxc}}[n_{i}]$ (10)
 which removes one-body self-interaction error and guarantees that the functional is exact for one-electron systems.
-Because evaluating the KIPZ functional is more expensive than the KI functional, it can be advantageous to evaluate the KIPZ Hamiltonian on the KI ground state.
+In contrast to KI, KIPZ does not share the same ground state density as the base functional and therefore mandates a full minimization of an orbital-density-dependent functional.
+Because of this, it can be advantageous to evaluate the KIPZ Hamiltonian on the KI ground state.
 This is referred to as perturbative KIPZ (pKIPZ).
-An important feature of Koopmans functionals is that they are not density functionals: the energy depends not only on the total density of the system but also on the densities of individual orbitals, making this an “orbital-density-dependent” functional theory.
-This comes with several important consequences.
-In the framework of DFT, the functional is invariant with respect to unitary rotations of the occupied manifold, and this means that the same set of orbitals minimize the total energy and diagonalize the Hamiltonian.
-However this is not the case for orbital-density dependent functionals, for which the “variational” orbitals minimize the energy while the “canonical” orbitals diagonalize the Hamiltonian constructed at the end of minimization procedure.
-These two sets of orbitals are related via a unitary transformation.
-However it is important to note that emergence of localized variational orbitals is crucial because it allows Koopmans functionals to correctly treat bulk systems [57].
+Because Koopmans functionals are orbital-density-dependent and have screening parameters that must be calculated ab initio, the procedure for computing a Koopmans band structure involves several steps as follows:
+  1. 1.
+    a ground-state DFT calculation to initialize the total density
+  2. 2.
+    a Wannierization of the DFT Kohn–Sham states in order to obtain maximally localized Wannier orbitals with which to initialize the variational orbitals
+  3. 3.
+    calculating the screening parameters $\{\alpha_{i}\}$ via finite differences or density functional perturbation theory [19, 17]
+  4. 4.
+    a final Koopmans calculation to minimize the energy and construct and diagonalize the Hamiltonian
+There are a few important technical details regarding this procedure.
+First, while any unitary rotation of the occupied manifold leaves the DFT energy unchanged, ODD functionals assign different energies to different sets of orbitals with the same total density.
+As a consequence, the minimization cannot rely on the usual self-consistent diagonalization procedure.
+Instead, the variational orbitals are obtained by directly minimizing the ODD energy using two nested loops, following the strategy introduced in ensemble DFT [77]: an inner loop optimizes the orbital rotations at fixed density, while an outer loop allows the density itself to relax.
+This conjugate-gradient procedure yields the self-consistent minimum of the functional (see Supplemental Material S1 [1] for more detail).
+Additionally, note that orbital occupations are not treated as variational parameters in this procedure.
+Since all systems considered here have an electronic gap, the occupied orbitals always have $f_{i}=1$ and the empty orbitals $f_{i}=0$, and only the orbitals themselves are optimized.
+Second, the emergence of variational orbitals — which are localized orbitals — is crucial because it allows Koopmans functionals to correctly treat bulk systems [85].
 The Koopmans correction is applied to each variational orbital, while the canonical orbitals are interpreted as Dyson orbitals and their eigenvalues as quasiparticle energies, which are shifted by some amount depending on the size of the Koopmans corrections applied to each of their constituent variational orbitals.
@@
 For systems with non-equivalent variational orbitals the Koopmans correction also affects inter- and intra-band distances, bandwidths, effective masses, etc..
-Because Koopmans functionals are orbital-density-dependent and have screening parameters that must be calculated ab initio, the procedure for computing a Koopmans band structure involves several steps as follows:
-  • a ground-state DFT calculation to initialize the total density
-  • a Wannierization of the DFT Kohn-Sham states in order to obtain maximally-localized Wannier orbitals with which to initialize the variational orbitals
-  • calculating the screening parameters $\{\alpha_{i}\}$ via finite differences or density functional perturbation theory [61, 62]
-  • a final Koopmans calculation to minimize the energy and construct and diagonalize the Hamiltonian
-The accuracy of Koopmans functionals has already been proven across a range of different materials [61, 63, 62, 64]. For more details about Koopmans functionals, we refer the reader to Ref. [65].
+This is discussed in more detail in Supplemental Material S2 [1].
+Third, the computational cost of Koopmans functional calculations is dominated by the calculations to obtain the screening parameters.
+This procedure is detailed in Supplemental Material S3 [1], including a discussion of how these calculations scale.
+In brief, the calculations to obtain the screening parameters amount to computing the energies of various charged defects.
+In this context, image charge corrections are especially important (see Supplemental Material S4 [1]).
+The accuracy of Koopmans functionals has already been demonstrated across a range of materials [19, 25, 17, 73]. For more details about Koopmans functionals, we refer the reader to Ref. [69].
 II.3 Band alignment with Koopmans functionals
 To calculate band alignment with Koopmans functionals, we must adapt the procedure described earlier in Section II.1.
-Here, we take advantage of a peculiar feature of the KI and pKIPZ functionals: for insulating systems where all the orbital occupancies are either 0 or 1, the ground state KI and pKIPZ energies and densities match those of the base functional.
-(This can be seen by setting $f_{i}$ to 0 or 1 in equation 9).
-We stress that this does not mean that the Koopmans correction has no effect, because the derivative of the correction is non-vanishing and thus the eigenvalues of the Koopmans functional will be different to those of the base functional.
-Nevertheless, this means that the ground-state density of Koopmans functionals match those of the base functional, and when performing band alignment the slab calculations only need to be performed at the level of DFT, because — being a ground-state property — the average electrostatic potential remains unchanged by the Koopmans correction [44, 66].
+Here, we take advantage of the aforementioned fact that ground state KI and pKIPZ energies and densities match those of the base functional.
+This means that when predicting band alignment with Koopmans functionals, the slab calculations only need to be performed at the level of DFT, because — being a ground-state property — the average electrostatic potential remains unchanged by the Koopmans correction [102, 99].
 This substantially reduces the cost of this framework for computing band alignment.
-Given this simplification, for Koopmans functionals equations 11 and 12 become:
+However, it is important to note that this comes with a caveat.
+Since the ground-state energies of the base functional and the corrected ones (KI@PBE or pKIPZ@PBE) are identical, this implies that neither PBE, KI@PBE, nor pKIPZ@PBE can reproduce the experimental phase stability of the reported TiO2 polymorphs.
+That being said, the third Koopmans functional KIPZ does not share this property and therefore modifies both the ground-state energy and density, and could therefore be used to address relative phase stability more reliably.
+However, applying KIPZ in the context of this work would require performing all the slab calculations fully at the KIPZ level, significantly increasing the computational cost.
+Proceeding with the KI and pKIPZ functionals, equations 11 and 12 become:
 $\mathrm{IP}=\Delta V^{\mathrm{DFT}}-(\varepsilon_{\mathrm{VBM}}^{\mathrm{DFT}}+\Delta\varepsilon_{\mathrm{VBM}})$ (11)
@@
 $\mathrm{EA}=\Delta V^{\mathrm{DFT}}-(\varepsilon_{\mathrm{CBM}}^{\mathrm{DFT}}+\Delta\varepsilon_{\mathrm{CBM}})$ (12)
-where $\Delta\varepsilon_{\mathrm{VBM}}$ and $\Delta\varepsilon_{\mathrm{VBM}}$ are the shifts in these band edges due to the Koopmans correction.
+where $\Delta\varepsilon_{\mathrm{VBM}}$ and $\Delta\varepsilon_{\mathrm{VBM}}$ are the shifts in the bulk band edges due to the Koopmans correction.
 Equivalently, one can calculate the electron affinity (EA) by simply subtracting the value of electronic gap calculated of Koopmans level from the IP:
@@
 Koopmans functionals calculations were performed using Quantum ESPRESSO via the koopmans package [69, 39, 38].
-The screening parameters were calculated via finite-differences, which necessitate the use of a supercell and image charge correction schemes to avoid spurious interactions between periodic images [77].
+The screening parameters were calculated via finite-differences, which necessitate the use of a supercell and image charge correction schemes to avoid spurious interactions between periodic images [76] (For more details, see Supplemental Material S4 [1]).
 For these calculations, we used a $2\times 2\times 2$ supercell for rutile and anatase and $2\times 1\times 2$ supercell for brookite.
@@
 These DFT calculations were also performed using Quantum ESPRESSO.
-The geometries of the slab structures were optimized in order to obtain an accurate physical picture of the surface/vacuum interface.
+The geometries of the slab structures were optimized using the PBE functional in order to obtain an accurate physical picture of the surface/vacuum interface.
 IV Results and discussion
@@
 The band gaps of rutile and brookite TiO2 are direct ($\Gamma\rightarrow\Gamma$), while that of anatase is indirect ($Z\rightarrow\Gamma$) [27, 123, 72].
-In all three cases, the KI correction shifts both the valence (downwards) and conduction bands (upwards), leading to the band gap widening relative to the PBE result.
-(The precise values of these shifts can be found in the Supplemental Material.) The resulting band gaps are presented in Table 1, alongside experimental values, as well as those calculated using hybrid functionals and GW.
+In all three cases, the KI correction opens the band gap, shifting the valence bands downwards and the conduction bands upwards.
+(The numerical shifts in the valence band maximum and conduction band minimum are provided in the Supplemental Material [1].) While the effect of the Koopmans correction may appear to be a rigid scissors shift, this is not the case; rather, it is an orbital-dependent correction whose effect on individual bands depend on the screening parameters and the orbital character.
+A full derivation of the KI potential and the conditions under which it reduces to a rigid shift is provided in Supplemental Material S2 [1].
+The resulting band gaps of the titania polymorphs are presented in Table 1, alongside experimental values, as well as those calculated using hybrid functionals and GW.
 When comparing against experimental band gaps it is important to account for zero-point renormalization — by subtracting the ZPR (as calculated ab initio) from experimental values one obtains a value that can be fairly compared against the computational results obtained for pristine crystal geometries.
@@
 rutile anatase brookite
-PBE $1.73$ $2.21$ $2.42$
+PBE $1.88$ $2.27$ $2.42$
 HSE06111[63]222[79] 3.39, 3.61 $3.60$ $3.30$
-G0W0@PBE11footnotemark: 1 $3.46$ $3.73$ $3.45$
-G0W0@HSE0622footnotemark: 2 $4.73$- --
-scQPGW@HSE0622footnotemark: 2 $5.18$- --
-KI@PBE $3.29$ $3.80$ $4.03$
-pKIPZ@PBE $3.38$ $3.90$ $4.14$
-exp -- ZPR $3.34-3.41$ $3.34-3.44$-
-exp333[93, 94, 95] $3.03-3.06$ $3.10-3.20$ $3.1-3.4$
-$-0.337,$
-ZPR444[96, 97, 98] $-0.349,$ $-0.238$-
-$-0.314$
+G0W0@PBE11footnotemark: 1333[54] $3.66,3.46$ $4.03,3.73$ $3.45$
+G0W0@HSE0622footnotemark: 2 $4.73$ - -
+GW0@PBE33footnotemark: 3 $4.23$ $4.60$ -
+QSGW444[14]22footnotemark: 2 $4.18,4.22$ - -
+QSGW̃44footnotemark: 4555[21] $3.73,3.88$ - -
+GW33footnotemark: 3 $4.84$ $5.28$ -
+KI@PBE $4.04$ $4.51$ $4.63$
+pKIPZ@PBE $4.19$ $4.64$ $4.78$
+exp - ZPR 3.14.13.1-4.1 3.43.93.4-3.9 -
+exp666[124, 119, 30, 107, 94, 90, 3, 33, 57, 58, 67, 95, 50] 2.83.82.8-3.8 3.23.73.2-3.7 3.13.53.1-3.5
+0.337,-0.337,
+ZPR777[80, 32, 118] 0.349,-0.349, 0.238-0.238 -
+0.314-0.314
 For rutile and anatase, the electronic band gaps calculated using KI@PBE are within 0.36 eV of experiment (when accounting for ZPR).
@@
 For brookite, no ZPR correction was found in the literature and thus comparison with experiment is not possible.
-The results with pKIPZ are very similar to that of KI (plots of pKIPZ@PBE band structures can be found in the Supplemental Material).
-Compared to the KI functional, pKIPZ reports slightly larger band gaps, especially in the case of brookite.
-Both KI and pKIPZ assign brookite the largest band gap of the three polymorphs, in contrast to HSE06 and G0W0@PBE which assign it the smallest.
-But without an accurate measurement of the ZPR in brookite and more precise experimental measurements of the band gap, it is difficult to determine which ordering is correct.
+[…]
+[…]
+Table 1 shows that the KI and KIPZ band gaps for rutile are in closer agreement with experiment than those for anatase and brookite, and that the differences among the three TiO2 polymorphs are more pronounced with KI and KIPZ than with HSE06 or $G_{0}W_{0}$.
+At present, no definitive explanation for this behavior can be offered.
+Given the wide spread in the reported experimental band gaps and the limited number of available data-points, drawing firm conclusions would be premature.
+There is no fundamental reason to expect differences between polymorphs to be inherently more pronounced when using Koopmans functionals than when using hybrid functionals or GW.
+Ongoing comprehensive benchmark studies are expected to clarify such trends more quantitatively.
+Finally, without an accurate measurement of the ZPR in brookite and more precise experimental measurements of the band gap, it is difficult to determine which ordering is correct.
 IV.2 Band alignment
@@
 IP EA IP EA IP EA
-PBE $7.22$ $5.55$ $7.11$ $4.90$ $7.17$ $4.78$
-HSE0600footnotetext: [33]11footnotemark: 1 $8.66$ $4.99$- -- -- --
-HSE06 (QM/MM)111[36] $7.83$- $8.3$- -- --
-G0W0@PBE11footnotemark: 1 $7.29$ $3.03$- -- -- --
-G0W0@HSE0611footnotemark: 1 $7.92$ $3.19$- -- -- --
-scQPGW@HSE0611footnotemark: 1 $8.77$ $3.59$- -- -- --
-KI@PBE $8.00$ $4.71$ $8.19$ $4.39$ $8.03$ $4.00$
-pKIPZ@PBE $7.83$ $4.44$ $8.00$ $4.10$ $7.84$ $3.71$
-exp -- ZPR $8.5$ $4.97-5.00$- -- -- --
-exp222[80, 99, 100] $8.20$ $5.14-5.17$ 7.96, 8.07, 8.20 -- -- --
-ZPR333[101] $-0.3$ $0.17$- -- -- --
+PBE $7.29$ $5.41$ $7.43$ $5.16$ $7.17$ $4.75$
+HSE0600footnotetext: [79]11footnotemark: 1 $8.66$ $4.99$ - - - -
+HSE06 (QM/MM)111[97] $7.83$ - $8.3$ - - -
+G0W0@PBE11footnotemark: 1 $7.29$ $3.03$ - - - -
+G0W0@HSE0611footnotemark: 1 $7.92$ $3.19$ - - - -
+scQPGW@HSE0611footnotemark: 1 $8.77$ $3.59$ - - - -
+KI@PBE $8.38$ $4.34$ $8.59$ $4.08$ $8.34$ $3.71$
+pKIPZ@PBE $8.14$ $3.95$ $8.33$ $3.69$ $8.08$ $3.30$
+exp - ZPR $8.5$ 4.975.004.97-5.00 - - - -
+exp222[108, 56, 109] $8.20$ 5.145.175.14-5.17 7.96, 8.07, 8.20 - - -
+ZPR333[26] 0.3-0.3 $0.17$ - - - -
 The ionization potentials and electron affinities of the $\mathrm{TiO_{2}}$ polymorphs, as calculated via KI and pKIPZ following the method described in Section II.3, are reported in Table 2 alongside values obtained from other higher levels of theory and experiment.
@@
 The EA, on the other hand, is predicted by KI@PBE within the range of values reported experimentally — a marked improvement upon the GW results, and matched only by HSE06.
-For anatase and brookite, no ZPR results for the IP and EA were found. The KI and pKIPZ result for the IP of anatase are fractionally lower than that given by HSE06 [36].
+For anatase and brookite, no ZPR results for the IP and EA were found.
+The KI and pKIPZ result for the IP of anatase are fractionally lower than that given by HSE06 [97].
 Figure 5: Band alignment of TiO2 polymorphs using KI and pKIPZ.
@@
 the VBM and CBM should not lie within the two dashed lines in Figure 5.
-According to our calculations, rutile TiO2 appears not to satisfy this condition, which may explain why — despite having a more desirable band gap than anatase and brookite — this polymorph is a slightly worse catalyst, especially compared to anatase [103].
-(Of course, an optimal band alignment does not guarantee photocatalytic success: anatase also exhibits higher surface activity and desirable excitonic properties [103, 104].) Brookite, on the other hand, shows some promise as a photocatalyst: the positions of the valence and the conduction bands seem to be favorable, but its fractionally larger band gap would limit the amount of photons that would be absorbed when subjected to sunlight.
+According to our calculations, rutile TiO2 appears to satisfy this condition, but only marginally, as its conduction-band minimum lies only slightly above the reduction potential.
+This borderline alignment may help explain why—despite having a more desirable band gap than anatase and brookite—rutile often performs as a less efficient photocatalyst, particularly relative to anatase[70].
+Brookite, on the other hand, shows some promise as a photocatalyst: the positions of the valence and the conduction bands seem to be favorable, but its fractionally larger band gap would limit the amount of photons that would be absorbed when subjected to sunlight.
 In this case, one might be able to engineer the band gap.
@@
 Still, with respect to the value of band gap and band edge positions, both KI and pKIPZ predict that anatase appears to be the most promising photocatalyst of the three.
+Of course, an optimal band alignment does not guarantee photocatalytic success: anatase also exhibits higher surface activity and desirable excitonic properties [70, 31].
+Indeed, excitonic binding energies and lifetimes vary significantly across the polymorphs.
+Rutile hosts weakly bound excitons (4–26 meV), with experiments reporting a 4 meV exciton [89] and GW calculations placing the lowest dipole-allowed exciton at 26 meV [55]; these small binding energies allow thermal dissociation, and charge carriers in rutile exhibit very short lifetimes [120].
+In contrast, anatase supports more strongly bound excitons, with GW calculations predicting a 160 meV exciton in the (001) plane [6], and a recent study including lattice distortions reporting exciton–polaron binding energies of 216 meV [23].
+Experiments also indicate that charge carriers in anatase are substantially longer-lived than in rutile [120].
+(To our knowledge, excitonic data for brookite are not available.) Importantly, these excitonic effects are not strong enough to appreciably shift the band-edge positions relative to the redox potentials and therefore do not influence the band-alignment results presented here.
 V Conclusions
@@
 VI Acknowledgments
-MS acknowledges support from EXAF (Excellence in Africa) Research Centre as part of the collaboration ”Embedded exact quantum dynamics for photocatalytic water splitting”.
+MS acknowledges support from EXAF (Excellence in Africa) Research Centre as part of the collaboration “Embedded exact quantum dynamics for photocatalytic water splitting”.
 This research was supported by the NCCR MARVEL, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 205602).